OpenStudy (anonymous):

MEDAL! What is the solution of the system of equations? 6x – 2y = 4 –9x + 3y = 12

OpenStudy (anonymous):

Given system of lines have no solution because these are parallel to each other. since a1/a2= b1=b2 i.e. a1/a2=6/(-9) = 2/(-3) b1=b2= -2/3 Hence a1/a2= b1=b2

OpenStudy (anonymous):

so no solution?